# More for Keys, Less for Values: Adaptive KV Cache Quantization

### v1

@@ 4.2 Experiment Results
Table 1 shows that across datasets and models, quantization error for the key matrices is consistently and substantially higher than that for the value matrices. For instance, in the MMLU dataset, Llama3.2-1B exhibits an error of $4.85\pm 1.04$ for $\mathbf{K}_{2}$ versus only $0.13\pm 0.10$ for $\mathbf{V}_{2}$. This trend persists for both the 3-bit and 4-bit settings, where increasing the bit-width reduces the error for both matrices, yet the keys still incur a notably higher error. Similar patterns are evident in the C4 and GSM8K datasets, highlighting that key matrices are inherently more sensitive to quantization noise. Furthermore, Figure 6 in the Appendix B illustrates the quantization error per attention head in each layer for the Llama3.1-8B model on C4, MMLU, and GSM8K, reinforcing our observation (see Appendix D for extended results). Table 2 shows accuracy on GSM8K under 1-shot and 8-shot prompting. Uniform K${}_{(2\text{bit})}$${}_{(2\text{bit})}$ settings degrade accuracy noticeably, especially in the longer 8-shot context. By contrast, K${}_{(4\text{bit})}$${}_{(2\text{bit})}$ nearly matches the unquantized baseline, confirming that providing higher precision to keys preserves model quality while still compressing values aggressively. The reversed assignment K${}_{(2\text{bit})}$${}_{(4\text{bit})}$ consistently underperforms, underscoring that allocating fewer bits to key matrices adversely affects accuracy.
5 Related Work
Model Quantization.
for large models such as LLMs Wan et al. (2024), which is why post-training quantization (PTQ) techniques are generally preferred for quantizing these models.
Quantization reduces the computational cost of neural network inference by lowering model bit-precision Han et al. (2016); Nagel et al. (2021). It can be divided into Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT). PTQ requires no re-training or labeled data Jacob et al. (2017); Nagel et al. (2019) and is preferred for large models like LLMs, whereas QAT demands fine-tuning with labeled data Nagel et al. (2021); Gholami et al. (2022) but scales poorly for such models Wan et al. (2024).
LLMs Quantization.
Researchers consider three post-training quantization settings for LLMs: weight-only, weight-activation, and KV cache quantization. Weight-only quantization focuses solely on weights Frantar et al. (2023a, b); Lin et al. (2024), while weight-activation quantization targets both weights and activations Dettmers et al. (2022); Xiao et al. (2023); Shao et al. (2024). Since longer inputs inflate the KV cache’s memory overhead Pope et al. (2023), KV cache quantization offers memory savings similar to weight-activation methods while maintaining performance near that of weight-only quantization Yue et al. (2024).
KV Cache Quantization.
Recent KV cache quantization methods reduce LLM memory and accelerate inference, falling into three categories: outlier redistribution, fixed-precision, and mixed-precision Li et al. (2025). Outliers in KV cache are addressed by redistributing them or applying smoothing transformations Xiao et al. (2023); Lin et al. (2024); Shao et al. (2024). Fixed-precision methods use a uniform bit-width Yao et al. (2022); Sheng et al. (2023); Liu et al. (2024) but can overlook token importance and outliers. In contrast, mixed-precision allocates higher precision to critical tokens and lower precision elsewhere Hooper et al. (2024); Yue et al. (2024); Li et al. (2024); Dong et al. (2024); Duanmu et al. (2024). Existing methods like SKVQ and QAQ rely on specific engineering solutions or predefined parameters, whereas KV-AdaQuant uses a principled, analysis-driven framework that leverages the intrinsic properties of key and value caches for a more robust and adaptable quantization approach.
@@ 6 Conclusion
In this work, we systematically analyze KV cache matrices from the perspectives of linear algebra and information theory to uncover distinct patterns in the information carried by these matrices. Building on the singular value distribution of the KV cache, we propose an information-aware quantization framework for KV cache compression. We evaluate the effectiveness of our approach on two widely used datasets, GSM8K, demonstrating its practical impact. While our current study focuses on the per-token quantization methods, the proposed framework is general and can be applied to any key-value quantization technique.
Limitations
Although our information-aware quantization framework is effective, there are aspects that warrant further consideration. First, our experiments utilize a maximum context length of 2K tokens. While this setting aligns with many practical inference scenarios, it may not fully capture the demands of tasks requiring larger context windows (e.g., 4K or 8K tokens). Extending our approach to support longer contexts could reveal additional challenges, such as increased quantization sensitivity and greater memory management overhead.
Ethical Considerations
Our research aims to reduce the memory footprint and computational overhead of LLM in inference time through KV cache quantization. Such advancements can potentially lower energy consumption and expand access to resource-constrained communities, contributing to more environmentally sustainable and inclusive deployment of language modeling based productions. However, the use of proposal KV cache compression method must be carefully monitored to ensure that accuracy trade-offs do not disproportionately affect critical applications such as healthcare, legal, and financial contexts.
Appendix A Quantization and Dequantization Error Bounds

### v2

@@ Results.
22footnotetext: Although Chain-of-Thought prompting is generally discouraged for reasoning models and can even impair their performance, we deliberately apply 1-shot and 5-shot CoT prompts here to examine the effect of longer input prompts on KV cache quantization.
Table 3: KV cache quantization comparison across GSM8K, EQ-Parseable, and CoQA datasets. The results are obtained on the HQQ quantization backend. Each cell shows accuracies for $\mathbf{K}_{i}\mathbf{V}_{x}$ and $\mathbf{K}_{x}\mathbf{V}_{i}$, where $x$ denotes the mean accuracy across quantization bit-widths $\{1,2,4,6,8\}$, with the percentage-point gap annotated above the comparison sign. See Appendix H for all settings.
GSM8K $\mathbf{K}_{1}\mathbf{V}_{x}$ vs. $\mathbf{K}_{x}\mathbf{V}_{1}$ $\mathbf{K}_{2}\mathbf{V}_{x}$ vs. $\mathbf{K}_{x}\mathbf{V}_{2}$ $\mathbf{K}_{4}\mathbf{V}_{x}$ vs. $\mathbf{K}_{x}\mathbf{V}_{4}$ $\mathbf{K}_{6}\mathbf{V}_{x}$ vs. $\mathbf{K}_{x}\mathbf{V}_{6}$ $\mathbf{K}_{8}\mathbf{V}_{x}$ vs. $\mathbf{K}_{x}\mathbf{V}_{8}$
Qwen3-0.6B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+3\%}}{{<}}\mathbf{0.03}$ $0.00\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{<}}\mathbf{0.16}$ $0.04\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{<}}\mathbf{0.17}$ $\mathbf{0.34}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.18$ $\mathbf{0.34}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.18$
Llama-3.2-1B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+5\%}}{{<}}\mathbf{0.05}$ $0.01\stackrel{{\scriptstyle\scriptscriptstyle+18\%}}{{<}}\mathbf{0.19}$ $\mathbf{0.27}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.20$ $\mathbf{0.28}\stackrel{{\scriptstyle\scriptscriptstyle+8\%}}{{>}}0.20$ $\mathbf{0.28}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.21$
Llama-3.2-3B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+19\%}}{{<}}\mathbf{0.19}$ $0.27\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{0.44}$ $\mathbf{0.57}\stackrel{{\scriptstyle\scriptscriptstyle+11\%}}{{>}}0.46$ $\mathbf{0.58}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.46$ $\mathbf{0.59}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}0.46$
Qwen3-4B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+41\%}}{{<}}\mathbf{0.41}$ $0.01\stackrel{{\scriptstyle\scriptscriptstyle+49\%}}{{<}}\mathbf{0.50}$ $\mathbf{0.80}\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{>}}0.51$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.51$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.51$
Llama-3.1-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{<}}\mathbf{0.31}$ $0.35\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{0.52}$ $\mathbf{0.70}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.54$ $\mathbf{0.70}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.54$ $\mathbf{0.70}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.54$
Qwen3-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+48\%}}{{<}}\mathbf{0.48}$ $0.08\stackrel{{\scriptstyle\scriptscriptstyle+46\%}}{{<}}\mathbf{0.54}$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.55$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{>}}0.55$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.55$
Qwen3-14B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+50\%}}{{<}}\mathbf{0.50}$ $0.26\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{<}}\mathbf{0.58}$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$
Qwen3-32B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+42\%}}{{<}}\mathbf{0.42}$ $0.25\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{<}}\mathbf{0.48}$ $\mathbf{0.73}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}0.50$ $\mathbf{0.71}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}0.50$ $\mathbf{0.72}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}0.51$
EQ-Parseable
Qwen3-0.6B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+10\%}}{{<}}\mathbf{10.06}$ $0.00\stackrel{{\scriptstyle\scriptscriptstyle+53\%}}{{<}}\mathbf{52.51}$ $51.23\stackrel{{\scriptstyle\scriptscriptstyle+2\%}}{{<}}\mathbf{52.87}$ $\mathbf{84.21}\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{>}}52.63$ $\mathbf{84.91}\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{>}}52.28$
Llama-3.2-1B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+34\%}}{{<}}\mathbf{33.68}$ $25.61\stackrel{{\scriptstyle\scriptscriptstyle+37\%}}{{<}}\mathbf{61.52}$ $\mathbf{87.25}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}65.38$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+24\%}}{{>}}66.32$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+25\%}}{{>}}66.08$
Llama-3.2-3B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+36\%}}{{<}}\mathbf{35.44}$ $68.19\stackrel{{\scriptstyle\scriptscriptstyle+5\%}}{{<}}\mathbf{73.10}$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}76.49$ $\mathbf{89.12}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}76.02$ $\mathbf{89.94}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}76.26$
Qwen3-4B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+54\%}}{{<}}\mathbf{47.60}$ $29.01\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{<}}\mathbf{58.01}$ $\mathbf{84.09}\stackrel{{\scriptstyle\scriptscriptstyle+28\%}}{{>}}59.88$ $\mathbf{85.96}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}59.06$ $\mathbf{85.38}\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{>}}59.88$
Llama-3.1-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+61\%}}{{<}}\mathbf{60.58}$ $\mathbf{77.89}\stackrel{{\scriptstyle\scriptscriptstyle+2\%}}{{>}}75.79$ $\mathbf{95.67}\stackrel{{\scriptstyle\scriptscriptstyle+19\%}}{{>}}77.19$ $\mathbf{96.96}\stackrel{{\scriptstyle\scriptscriptstyle+20\%}}{{>}}77.31$ $\mathbf{97.43}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}77.08$
Qwen3-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+62\%}}{{<}}\mathbf{58.48}$ $54.39\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{70.41}$ $\mathbf{95.32}\stackrel{{\scriptstyle\scriptscriptstyle+25\%}}{{>}}71.35$ $\mathbf{96.26}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}70.76$ $\mathbf{96.14}\stackrel{{\scriptstyle\scriptscriptstyle+26\%}}{{>}}71.11$
Qwen3-32B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+60\%}}{{<}}\mathbf{47.37}$ $56.14\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{<}}\mathbf{61.75}$ $\mathbf{80.12}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}62.46$ $\mathbf{80.47}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}62.34$ $\mathbf{79.77}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}62.57$
CoQA
Qwen3-0.6B $0.20\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{<}}\mathbf{0.44}$ $0.32\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{<}}\mathbf{0.52}$ $\mathbf{0.65}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.53$ $\mathbf{0.69}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}0.53$ $\mathbf{0.69}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}0.53$
Llama-3.2-1B $0.22\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{<}}\mathbf{0.42}$ $0.48\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{<}}\mathbf{0.57}$ $\mathbf{0.67}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}0.58$ $\mathbf{0.68}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.58$ $\mathbf{0.68}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.58$
Llama-3.2-3B $0.21\stackrel{{\scriptstyle\scriptscriptstyle+50\%}}{{<}}\mathbf{0.60}$ $\mathbf{0.73}\stackrel{{\scriptstyle\scriptscriptstyle+6\%}}{{>}}0.68$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.68$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.68$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+15\%}}{{>}}0.67$
Qwen3-4B $0.34\stackrel{{\scriptstyle\scriptscriptstyle+34\%}}{{<}}\mathbf{0.62}$ $0.68\stackrel{{\scriptstyle\scriptscriptstyle+3\%}}{{<}}\mathbf{0.70}$ $\mathbf{0.81}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.71$ $\mathbf{0.81}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.71$ $\mathbf{0.81}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}0.70$
Llama-3.1-8B $0.21\stackrel{{\scriptstyle\scriptscriptstyle+50\%}}{{<}}\mathbf{0.60}$ $\mathbf{0.72}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.66$ $\mathbf{0.78}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.67$ $\mathbf{0.78}\stackrel{{\scriptstyle\scriptscriptstyle+15\%}}{{>}}0.67$ $\mathbf{0.78}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.67$
Qwen3-8B $0.38\stackrel{{\scriptstyle\scriptscriptstyle+37\%}}{{<}}\mathbf{0.68}$ $\mathbf{0.73}\stackrel{{\scriptstyle\scriptscriptstyle+1\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$
Qwen3-14B $0.32\stackrel{{\scriptstyle\scriptscriptstyle+45\%}}{{<}}\mathbf{0.69}$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+8\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$
Qwen3-32B $0.32\stackrel{{\scriptstyle\scriptscriptstyle+44\%}}{{<}}\mathbf{0.68}$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+9\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.73$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.73$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.73$
@@ 5 Conclusion
Large scale language models increasingly spend the majority of their inference footprint on the KV cache, making cache compression a decisive factor in practical deployment. This work supplied a principled answer to the bit-allocation problem by unifying theory and experiment.
  • Key–Value Norm Disparity. We proved that Key weights exhibit larger Frobenius norms than Value weights throughout training, implying a higher information density in the key path.
  • Key-Driven Quantization. Leveraging this norm gap, a second theorem shows that assigning the finer bit-width to keys and the coarser bit-width to values strictly minimises quantization error, thereby preserving downstream accuracy.
  • Empirical Validation. Experiments across nine model families, three generative benchmarks, and two hardware-aligned back-ends confirm the theory: a $4\!\times\!2$ (keys / values) split consistently recovers $\geq 94\%$ of full-precision accuracy while halving KV-cache memory.
These findings elevate bit allocation from a hyper-parameter tuned by grid search to a design principle derived from model geometry, offering immediate guidance for practitioners and a foundation for hardware–algorithm co-design. We used a context length of 2K tokens, a common setting, but will use longer contexts in future.
Appendix
Appendix A Dynamics of the Norms of Key and Value Weight Matrices
A.1 Preliminary: Dimensions
$\begin{array}[]{@{}ll@{}}c&\text{sequence length (context size)},\\[4.0pt] d_{m}&\text{embedding dimension},\\[4.0pt] d_{v}=d_{k}=d_{m}/n_{h}=d_{m}&\text{for a single head ($n_{h}=1$)},\\[4.0pt] n_{h}&\text{number of attention heads}.\end{array}$
A.2 Symbols
$\begin{array}[]{@{}ll@{}}X&\text{input matrix }(c\times d_{m}),\\ X^{O}&\text{output matrix }(c\times d_{m}),\\ W^{V}&\text{value weights }(d_{m}\times d_{m}),\\ W^{K}&\text{key weights }(d_{m}\times d_{m}),\\ W^{Q}&\text{query weights }(d_{m}\times d_{m}),\\ K=XW^{K}&(c\times d_{m}),\\ Q=XW^{Q}&(c\times d_{m}),\\ V=XW^{V}&(c\times d_{m}),\\ S=QK^{\top}/\sqrt{d_{m}}&(c\times c),\\ A=\operatorname{softmax}(S)&(c\times c),\\ \mathcal{L}&\text{loss function.}\end{array}$
Appendix B Norm Dynamics
In this appendix, we compare the long–time behavior of the Frobenius norms of the key ($W^{K}$) and value ($W^{V}$) weight matrices of a single–head self–attention layer trained with stochastic gradient descent (SGD). We show that, under standard isotropic assumptions, $\|W^{K}\|_{F}$ grows faster than $\|W^{V}\|_{F}$.
B.1 Update rule
At update step $i$ the weights follow
$W^{m}_{i+1}\;=\;W^{m}_{i}-\eta\,\frac{\partial\mathcal{L}}{\partial W^{m}_{i}},\qquad m\in\{K,V\},$ (2)
with learning rate $\eta$. Squaring the Frobenius norm of both sides gives
$\|W^{m}_{i+1}\|_{F}^{2}$ $=\|W^{m}_{i}\|_{F}^{2}+\eta^{2}\Bigl{\|}\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\|}_{F}^{2}-2\eta\,\Bigl{\langle}W^{m}_{i},\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\rangle}_{F},$ (3)
where
$\|A\|_{F}$ $\equiv\sqrt{\operatorname{tr}(A^{\top}A)},$ (4)
$\langle A,B\rangle_{F}$ $\equiv\operatorname{tr}(A^{\top}B).$ (5)
B.2 Expectation over mini-batches
Taking an expectation over mini-batches yields
$\Delta\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}$ $\equiv\mathbb{E}\bigl{[}\|W^{m}_{i+1}\|_{F}^{2}\bigr{]}-\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}$ (6)
[…]
In high–dimensional weight space the gradient is almost orthogonal to the current weights, making the second term negligible. Hence,
$\Delta\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}\;\approx\;\eta^{2}\,\mathbb{E}\Bigl{[}\Bigl{\|}\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\|}_{F}^{2}\Bigr{]}.$ (8)
B.3 Initialization
Throughout we assume Xavier initialization: each entry of $W^{K}$ and $W^{V}$ is drawn i.i.d. from a zero–mean distribution whose variance preserves the input scale [36].
B.4 Gradients
The attention outputs are
$X^{O}\;=\;A\,V\,W^{O},$ (9)
where $A=\operatorname{softmax}\!\bigl{(}QK^{\top}/\sqrt{d_{m}}\bigr{)}$ and $V=XW^{V}$. Differentials give
$d\mathcal{L}(X^{O})$ $=\operatorname{tr}\!\bigl{[}(\partial\mathcal{L}/\partial X^{O})^{\!\top}\,AX\,dW^{V}W^{O}\bigr{]}=\operatorname{tr}\!\bigl{[}(AX)^{\top}\,\frac{\partial\mathcal{L}}{\partial X^{O}}\,(W^{O})^{\top}\,dW^{V}\bigr{]},$ (10)
so that
$\frac{\partial\mathcal{L}}{\partial W^{V}}=(AX)^{\top}\,\frac{\partial\mathcal{L}}{\partial X^{O}}\,(W^{O})^{\top}.$ (11)
Similarly, writing $S=QK^{\top}/\sqrt{d_{m}}$,
$d\mathcal{L}(S)$ $=\operatorname{tr}\!\bigl{[}(\partial\mathcal{L}/\partial S)^{\!\top}\,dS\bigr{]}=\frac{1}{\sqrt{d_{m}}}\,\operatorname{tr}\!\bigl{[}X^{\top}\,(\partial\mathcal{L}/\partial S)^{\!\top}\,Q\,dW^{K}\bigr{]},$ (12)
yielding
$\frac{\partial\mathcal{L}}{\partial W^{K}}=\frac{1}{\sqrt{d_{m}}}\,X^{\top}\,(\partial\mathcal{L}/\partial S)^{\!\top}\,Q.$ (13)
B.5 Expected gradient norms
Squaring the Frobenius norm of (11) and taking an expectation under the isotropic-input assumption,
$\mathbb{E}\!\bigl{[}\|\partial\mathcal{L}/\partial W^{V}\|^{2}_{F}\bigr{]}$ $=c\,\sigma_{x}^{2}\,\sigma_{o}^{2},$ (14)
where $\mathbb{E}[X^{\top}X]=c\sigma_{x}^{2}I$ and $\mathbb{E}\Bigl{[}\Bigl{(}A^{\top}\tfrac{\partial\mathcal{L}}{\partial X^{O}}(W^{O})^{\top}\Bigr{)}\Bigl{(}A^{\top}\tfrac{\partial\mathcal{L}}{\partial X^{O}}(W^{O})^{\top}\Bigr{)}^{\top}\Bigr{]}=\sigma_{o}^{2}I$.
For the key weights,
$\mathbb{E}\!\bigl{[}\|\partial\mathcal{L}/\partial W^{K}\|^{2}_{F}\bigr{]}=\frac{\sigma_{x}^{2}\,\sigma_{s}^{2}}{d_{m}}\,\|Q\|_{F}^{2},$ (15)
with $\sigma_{s}^{2}$ the entry-wise variance of $\partial\mathcal{L}/\partial S$.
B.6 Asymptotic comparison
Substituting these expectations into (8) shows that
$\mathbb{E}\!\bigl{[}\|W^{K}\|^{2}_{F}\bigr{]}\text{grows as}\;\eta^{2}\,\frac{\sigma_{x}^{2}\sigma_{s}^{2}}{d_{m}}\,\|Q\|_{F}^{2},\quad\text{whereas}\quad\mathbb{E}\!\bigl{[}\|W^{V}\|^{2}_{F}\bigr{]}\;\text{grows as}\;\eta^{2}\,c\sigma_{x}^{2}\sigma_{o}^{2}.$ (16)
Because $\|Q\|_{F}^{2}$ itself increases during training, $\|W^{K}\|_{F}$ eventually dominates:
$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}\bigr{]}\;>\;\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}\bigr{]}.$ (17)
As shown in Equation 17, the expectation of the Frobenius norm of $W^{k}$ is greater than that of $W^{v}$. [^3] While this analysis assumes a single-head, decoder-only transformer for clarity, the extension to multi-head attention is direct. The same conclusion also holds for group-query and multi-query attention.
Appendix C Quantization and Dequantization Error Bounds